OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[3x \sqrt{2}\times \sqrt{14x}\]

OpenStudy (anonymous):

6(7^1/2)x^3/2

OpenStudy (anonymous):

@zzr0ck3r can you help me with this

OpenStudy (zzr0ck3r):

3*x*sqrt(2)*sqrt(14x) = 3*x*sqrt(2*14x) = 3*x*sqrt(28x) = 3*x*sqrt(4*6*x) = 3*x*sqrt(4)*sqrt(6x) = 3*x*2*sqrt(6x) = 6*x*sqrt(6x)

OpenStudy (zzr0ck3r):

this is all because sqrt(a*b) = sqrt(a) * sqrt(b)

OpenStudy (anonymous):

Thats not one of the answer choices

OpenStudy (zzr0ck3r):

what are they?

OpenStudy (zzr0ck3r):

o wait

OpenStudy (anonymous):

6xsqrt7x 7xsqrt6x 12xsqrtx 21xsqrt2x

OpenStudy (anonymous):

In place of 6x,there will be 7x in ur answer,zzr0ck3r.

OpenStudy (zzr0ck3r):

yeah what he said^

OpenStudy (anonymous):

So it would be the first answer choice

OpenStudy (zzr0ck3r):

28 = 4*7 not 4*6 lol sorry

OpenStudy (zzr0ck3r):

yes, do you understand why?

OpenStudy (anonymous):

Yea i saw from your steps can you help me with one more

OpenStudy (zzr0ck3r):

close this and post another, so other can help @Swati19 was right, but some people don't like fractions in the exponent if not needed.

OpenStudy (zzr0ck3r):

or just post it here because you are already typing:)

OpenStudy (anonymous):

\[5\sqrt{3( 2\sqrt{3})}+6\sqrt{19}\]

OpenStudy (zzr0ck3r):

sec, phone. write them as exponents and it will help

OpenStudy (zzr0ck3r):

\[5(3*2*3^{\frac{1}{2}})^{\frac{1}{2}}+6(19)^{\frac{1}{2}}=5(6^{\frac{1}{2}}*3^{\frac{1}{4}})+6(19)^{\frac{1}{2}}\]

OpenStudy (anonymous):

30+30sqrt57???

OpenStudy (zzr0ck3r):

what are all the choices, there are so many ways to change this...

OpenStudy (anonymous):

30+30sqrt57 90+30sqrt57 10sqrt3 + 30sqrt19 10sqrt3 + 30sqrt57 @zzr0ck3r

OpenStudy (zzr0ck3r):

30+30sqrt57 = 256.49 90+30sqrt57 = 316.495 10sqrt3 + 30sqrt19 = 148.087 10sqrt3 + 30sqrt57 = 243.81 but 5sqrt(6*sqrt(3))+6sqrt(19) = 42.27 are you sure you wrote the question right

OpenStudy (anonymous):

Yea  @zzr0ck3r